OpenStudy (anonymous):

Help! need step by step to understand 4x/9 + x-10/9 t^2+6t/t-1 - 7/t-1 3/x + 13/x-11 5/4x^2y - y/ 14xz x^2+2x/12x+54 - 3x/8x+36 2/x^2+8x+15 + 1/x^2+11x+30

OpenStudy (phi):

you should use parentheses x - 10/9 means \( x - \frac{10}{9}\) but you probably mean (x-10)/9 which means \( \frac{x-10}{9} \)

OpenStudy (phi):

if two fractions have the same denominator, you add the tops, and the bottom stays the same example: \[ \frac{1}{4} + \frac{2}{4} = \frac{1+2}{4} = \frac{3}{4} \] this rule works for letters Example: \[ \frac{x}{a} + \frac{2y}{a} = \frac{x+2y}{a} \]

OpenStudy (anonymous):

for that i got 5x+10/9?

OpenStudy (phi):

first, use parens 2nd, how did the -10 change to +10 ?

OpenStudy (anonymous):

i meant 5x-10/9

OpenStudy (phi):

you meant (5x-10)/9 both the 5x and the -10 are being divided by 9

OpenStudy (anonymous):

yes. so thats my final answer since its just simplifying right?

OpenStudy (phi):

what did you get for the 2nd question ?

OpenStudy (anonymous):

(t+7)

OpenStudy (phi):

\[ \frac{t^2+6t-7}{t-1} \] but you can factor the top into (t+7)(t-1) so this is \[ \frac{(t+7)(t-1)}{t-1} = t+7 , t≠1\]

OpenStudy (phi):

for the 3rd question you need a common denominator. in this case x(x-11) is the smallest common denominator

OpenStudy (anonymous):

okay. for the third i got (16x-33)/x^2-11x

OpenStudy (phi):

yes, but I would type it (16x-33)/(x^2-11x) because that whole mess (x^2-11x) is in the bottom It looks like you know what you are doing.

OpenStudy (anonymous):

the only real problem im having is number 6. can you help me with that?

OpenStudy (phi):

the 4th question also needs a common denominator try 4*14*x^2*y*z

OpenStudy (anonymous):

can you help with 6? @phi

OpenStudy (phi):

\[ \frac{2}{x^2+8x+15}+ \frac{1}{x^2+11x+30}\] I assume you know how to factor these? if not see http://www.khanacademy.org/math/trigonometry/polynomial_and_rational/quad_factoring/v/factoring-quadratic-expressions you get \[ \frac{2}{(x+3)(x+5)}+ \frac{1}{(x+5)(x+6)}\] the common denominator will be (x+3)(x+5)(x+6) that means multiply the first fraction by (x+6)/(x+6) and multiply the 2nd fraction by (x+3)/(x+3) can you finish ?